# CU-JADE: A Method for Traversing Extinction Jumps along the Line-of-Sight

### v1

@@ 5.1 All Sky Dust Map
This method does not require prior assumptions of the number of jump points. Indeed, due to the complex distribution of gas on the Galactic plane, it is difficult to preset how many extinction components there are along a certain LoS. For instance, the $D$-$A$ curve is not merely a simple step function, but involves combinations of different scenarios (see details in Appendix B). Our method can traverse the jump points along the LoS in a certain region. This can be applied to detect components with significant extinction across the full sky, effectively generating a 3D extinction map for whole sky (see Figures 4 and 5 for the sky map and top-down views).
@@
To map all LoS characteristics across the celestial sphere, we apply the Cylindrical Equal Area Rectangular (CAR) projection, which projects the sphere onto a rectangular plane. We set the sampling size (represent spatial resolution) $L_{0}=1^{\circ}$ and sampling interval (grid size) $L=30^{\prime}$ for a guide map. Setting aside the notable variations in star density across high Galactic latitudes and the Galactic plane, we generate a uniformly sampled planar diagram with a $30^{\prime}$ grid spacing and $30^{\prime}$ overlap to enhance the spatial correlation. This standardized format facilitates easy comparison with other maps. However, unlike traditional extinction maps, the distribution of extinction jumps in our map is discrete in the dimension of distance. For the sky map, we perform integration across various distance slices, while for face-on map, we employ a two-dimensional histogram with extinction magnitude as the weighting factor.
@@ 5.1.1 A Dust Map based on SHEDR3
We utilized the dataset from SHEDR3 (by Mode I, $C=2$), providing measurements of extinction $A_{V}$ and distance, with the maximum star distance capped at 5 kpc. The integration of this star table with infrared data aids in detecting more distant components with deeper extinction. The magnitudes of all identified jump points adhere to an exponential distribution. To improve detection accuracy and focus on more significant jumps, we retain jumps with $\Delta A_{V}$ exceeding 0.15 mag 111 A sample of this catalog with the description of the columns can be found in Table 2. The full machine-readable table of extinction components is presented in 10.57760/sciencedb.24650, from which the 3D extinction maps and distances of MCs can be derived.. This threshold is set by considering three key reasons: (1) 0.15 mag represents a typical precision value for extinction in the $V$ band from SHEDR3 at $G=17$ mag (Anders et al., 2022); (2) Section 4.1’s simulations demonstrate the algorithm’s effectiveness in detecting jump points above this threshold (assuming $\sigma\sim 0.2-0.5$ mag); (3) The most important is, at this threshold, molecular hydrogen might be experiencing a phase transition based on the conversion between the extinction magnitude and the total column density of all gas. Here, we apply the empirical relationship between HI and the reddening, given by $N_{\rm HI}/E(B-V)=8.8\times 10^{21}{\rm cm}^{-2}{\rm mag}^{-1}$ (Lenz et al., 2017), with an extinction coefficient $R_{V}=3.1$ (Cardelli et al., 1989; Schlafly & Finkbeiner, 2011). This implies that the extinction contributed by HI is $A_{V}({\rm HI})=0.35\times 10^{-21}N_{\rm HI}$. And the total extinction $A_{V}$ reflects a combination of atomic gas and molecular hydrogen, which is $N_{\rm H}=N({\rm HI})+2N({\rm H}_{2})\approx 1.9\times 10^{21}{\rm cm}^{-2}A_{V}$ (Bohlin et al., 1978). Based on empirical relationships, the column density range where molecular hydrogen appears is $log~{}N_{\rm H}({\rm cm}^{-2})\approx 20.1-20.8$ (Gillmon et al., 2006; Bellomi et al., 2020), corresponding to an extinction value of $A_{V}({\rm mag})\approx 0.06-0.3$, with 0.15 mag being a mid-range value. Concurrently, Dame et al. (2001) suggests an extinction threshold for CO of $\gtrsim 0.4$ mag, indicating that some CO-dark molecular gas might appear below $\sim 0.4$ mag (Kalberla et al., 2020; Skalidis et al., 2024). We expect that a higher sensitivity survey may detect weak CO emission in some CO-dark regions.
@@
$\mu^{\mathrm{fore}}$: The mean extinction in the pre-jump distance bin.
$\mu^{\mathrm{back}}$: The mean extinction in the post-jump distance bin.
$\sigma^{\mathrm{fore}}$: The standard deviation of extinction in the pre-jump distance bin.
$\sigma^{\mathrm{back}}$: The standard deviation of extinction in the post-jump distance bin.
@@
In Figure 4, the Mollweide projection plot of the sky map illustrates the distances associated with the most significant extinction jumps across all LoS. The diagram shows some notable nearby regions that likely match those identified MCs by the single-extinction-jump model (Yan et al., 2019a; Zucker et al., 2019). The advantages of our model are that (1) the model with multi-extinction-jump can well trace all extinction structures with considerable $\Delta A_{\lambda}$ and (2) all the detected dust components are unaffected by different modeling process, i.e., the “ramps” in the $D$-$A$ map (Zucker et al., 2019) and the overlapping between multiple gas layers in distance measurements (i.e., multiple jump features along LoS, see Zhang et al., 2024). For the majority of the LoS, particularly at high Galactic latitudes, MCs are usually isolated and correspond to the strongest extinction components, facilitating the identification of their associated extinction structures and distances (Sun et al., 2024a). In contrast, the Galactic plane often features multiple molecular gas layers associated with various extinction components, complicating the precise determination of their distances. Given the correlation between MC column density and extinction, we prioritize the analysis of more significant extinction components (see Sections 5.2.1 and 5.2.2).
@@
Recent studies utilizing Gaia parallaxes have shed new light on the structure of the solar neighborhood. For example, some nearest star-forming regions are found to be located near the surface of the Local bubble (e.g., Taurus, Ophiuchus, Lupus, Chamaeleon, and Corona Australis; see details in Zucker et al., 2022). Our results also show prominent extinction features toward these regions within 300 pc (see thick blue regions in Figure 4). Some intriguing extinction structures are also revealed by our map (see light yellow bands for kpc-scale linear structures in Figure 5), e.g., the spur-like structure between the Sagittarius and Local arms (“Split” in Lallement et al., 2019), the Radcliffe Wave (Alves et al., 2020), and the Cepheus spur (Pantaleoni González et al., 2021), etc. Besides, the large cavity in $120^{\circ}\lesssim l\lesssim 135^{\circ}$ (green ellipse in Figure 5, Vergely et al., 2022) is also clear in our map. We also find more oval cavity structures that dust is scarce in $200^{\circ}\lesssim l\lesssim 210^{\circ}$, $280^{\circ}\lesssim l\lesssim 300^{\circ}$ (see red ellipses in Figure 5). We call them the Mon-Gem superbubble and the Carina superbubble, respectively. While conventional extinction mapping obscures these features due to severe line-of-sight extinction and artificial structures, our analysis more clearly reveals their kpc-scale morphology. These giant bubbles are reminiscent of the vast cavities observed across NGC 628 (the Phantom Voids; Watkins et al., 2023; Barnes et al., 2023), likely reflecting the global impact of intense past star formation activity on the gas distribution and dynamics of the Galactic disk. Future multi-wavelength (e.g., X-ray, infrared) observations will be essential for further verification and confirm their formation mechanisms.
Figure 6: Face-on maps from different star catalogs/techniques. (a) and (b) present Polar coordinate dust maps of ZGR23. (a) is produced by our method, and (b) is generated using a simple spatial binning technique as outlined by Zhang et al. (2023). (c) is reproduced in Cartesian coordinates from the map in Sect 5.1.1. (d) is from Vergely et al. (2022), using a hierarchical inversion technique and two merged star catalogs.
@@ 5.1.2 Qualitative Comparison of Different Dust Maps
Figure 6 illustrates the following results: Panels (a) and (b) both utilize the ZGR23 catalog in polar coordinates, with (a) represents our proposed method and (b) using a simple spatial binning-based mapping technique. Notably, our method effectively suppresses finger-like artifacts and recovers structures in distant and low-extinction regions. Moreover, the reconstruction of the Perseus Arm in the second and third Galactic quadrants is particularly robust. While both methods maintain consistency in large-scale structures near the Sun, our approach demonstrates significant improvements in resolving finer details in complex environments.
@@
Panel (c) in Figure 6 presents a Cartesian projections of Figure 5, while (d) display hierarchical-inversion extinction map from Vergely et al. (2022). Both methodologies share three key traits: they utilize model-independent approaches, rely on the Gaia EDR3 astrometric backbone, and incorporate extensive optical/infrared survey data. Their top-down views align well in the solar neighborhood and inner Galaxy, but notable discrepancies emerge in some specific features. In essence, panel (c) shows a cleaner, more visualization by effectively mitigating “fingers-of-God” artifacts. Panel (d), while offering higher resolution, displays discrepancies such as the Scutum-Centaurus Arm appearing more distant and dispersed, and MCs like the one at (1.7, -3) kpc appearing farther. Indeed, panel (c) from our work exhibits significantly greater spiral arm coherence than panel (d), especially in regions towards both the inner ($1.6\pm 0.2$ kpc and $2.6\pm 0.2$ kpc toward the Galactic center; Nogueras-Lara et al., 2021) and outer Galaxy (e.g., Sun et al., 2024b).
@@
Panels (a) and (c) in Figure 6 utilize the same methodology but rely on different catalogs. Their face-on views reveal systematic differences in their extinction estimates. While panel (a) results in a more compact distribution of dust clouds with reduced granularity, panel (c) more effectively resolves faint structures and exhibits prominent continuity in spiral arm features. The most notable differences between the two panels occur within the inner Galaxy, particularly towards the Galactic center, where panel (c) displays fewer elongated patterns in the face-on map but a sharper spiral feature.
@@
Previous studies (e.g., Green et al., 2019; Edenhofer et al., 2024) have conducted comparative analyses of various 3D extinction maps. However, significant methodological differences in dust mapping techniques and extinction correction approaches hinder direct comparisons between these studies. To establish definitive conclusions, controlled analyses using identical methods and stellar catalogs are essential. While variations exist in distance coverage and spatial resolution across different maps, these factors primarily affect small-scale structures rather than the overall large-scale morphology of the solar neighborhood. In summary, the CU-JADE method enables detection of MC distances even from weak extinction jumps, while simultaneously improving 3D extinction mapping through reduction of fingers-of-God artifacts.
@@ 5.2.1 Cepheus Region
The MCs in the Cepheus region with different distances were roughly divided into three groups based on previous studies (Grenier et al., 1989; Yonekura et al., 1997): the nearby (200 pc $\lesssim d\lesssim$ 400 pc; ${\rm-9~{}km~{}s}^{-1}\lesssim v\lesssim$ 9 ${\rm km~{}s}^{-1}$), the distant (600 pc $\lesssim d\lesssim$ 1000 pc; ${\rm-26~{}km~{}s}^{-1}\lesssim v\lesssim$ 2 ${\rm km~{}s}^{-1}$), and those in the Perseus Arm ($d\gtrsim$ 3 kpc; ${\rm-37~{}km~{}s}^{-1}\lesssim v\lesssim$ ${\rm-30~{}km~{}s}^{-1}$). The distant group might overlap with the Cepheus Flare at high Galactic latitudes ($b\sim 13^{\circ}$), leading to large uncertainty and debate over the distances in this area. Zucker et al. (2019) also measured a group of distances at $\sim 950$ pc and 350350~{}350350 pc toward the region using a single-extinction-jump model.
@@
Figure 7: CfA $\rm{}^{12}CO$ and extinction map toward the Cepheus region. Panel (a): Depicts the extinction between 200 and 500 pc using a colormap, while contours show the smoothed integrated intensity of $\rm{}^{12}CO$ emission within the LSR velocity range of -6 to 15 $\rm km~{}s^{-1}$. Panel (b): Similar to panel (a), but focuses on the distance range in (500, 1500) pc and LSR velocities in (-20, -6) $\rm km~{}s^{-1}$. Panel (c): Similar to panel (a), but presents data for distances in (1500, 3000) pc and LSR velocity in (-45, -20) $\rm km~{}s^{-1}$. Red circles indicate regions where extinction jumps were detected in right panels (d), (e), and (f). Vertical lines represent the distances of detected extinction jumps with uncertainties in $D$-$A$ plots. Grey dots are raw stellar data points, binned every 50 pc with black lines. Blue text and horizontal lines denote the mean extinction value after accounting for extinction jumps. Orange curves are binned radial velocities of stars from the Gaia DR3.
@@
Combined with CO data, our results give a more comprehensive view on this area. In Figure 7, we presented the zoom-in region between $l\in[100^{\circ},120^{\circ}]$ and $b\in[0^{\circ},20^{\circ}]$ from the all-sky map in Figure 4. Generally, MCs in Cepheus exhibit good correlations between the velocity and the corresponding distance, as well as the total gas column density and extinction (see panels (a), (b), and (c) in Figure 7). Based on the dust map in the Figure 7 (b), we have well determined some extended extinction structures from high Galactic latitudes to Galactic plane (see structures at $500-1500$ pc). These extinction structures matches well with the molecular gas in distant groups (black contours in panel (b) of Figure 7). This large-scale structure may extend to the junction of Cepheus and Cassiopeia (see Figure 4), where the star burst cluster M120.1+3.0 is located at $\sim 1$ kpc (Sun et al., 2023). For MCs in the most negative velocities ($v\lesssim-20$ ${\rm km~{}s}^{-1}$), the corresponding extinction structures are primarily located in the Galactic plane and extend to the $b\sim 5^{\circ}$ (see circle f in panel (c) of Figure 7).
@@
In this direction, the nearby MCs at 200–500 pc (contours and colormap in panel (a) of Figure 7) are widely and non-uniformly distributed throughout the whole region. Studies suggest a concentration of young stars aged 1–5 Myr at distances of 330–370 pc (Szilágyi et al., 2021), indicating that these young stellar groups may be associated with the surrounding molecular gas at this distance. The nearby MCs with ${\rm-6~{}km~{}s}^{-1}\lesssim v\lesssim$ 15 ${\rm km~{}s}^{-1}$ are roughly ring-shaped, while the distant MCs (500–1500 pc; see circle e in panel (b) of Figure 7) are interspersed among them. These distant MCs near 1 kpc cause more significant extinction at $l\sim 109^{\circ},b\sim 13^{\circ}$ than structures at $\sim 300$ pc (see panel (e) of Figure 7). Our results show that there are two or even more gas layers toward the area near $l\sim 107^{\circ}-111^{\circ}$and $b\sim 13^{\circ}$ in the Cepheus Flare. Indeed, the Cephues Flare region exhibits multiple star-formaion episodes (about 7–30 Myr; Wang et al., 2025) that appear to have significantly reshaped the gas structure and spatial distribution within 500 pc.
@@
The $D$-$A_{G}$ diagrams (by Mode II and $C$=2) of three directions (labeled d, e, and f in Figure 7) are collected from star samples within the circles. Combined with the left side-by-side plots, we further confirmed the close physical connection between CO clouds with different LSR velocities and the corresponding extinction structures in gas layers at different distances. At the same time, we plot the stellar radial velocities (RV) after binning along the distance. The RV of the stars only roughly conforms to the trend of the rotation curve. However, we find that the dispersion and the uncertainty of the RV of stars are very large (see orange lines in $D$-$A_{G}$ diagram of Figure 7). Particular caution should be taken when using stellar RV in small velocity ranges and distance intervals to study the kinematics of their structures. On the contrary, the RV of young star might be useful for describing the kinematics of the young stellar objects because of their smaller kinematical ages (Zhu et al., 2024).
@@
Figure 8: Overlay of the identified MWISP $\rm{}^{13}CO$ cloud structures (white contours) within 1.1 kpc (clouds in the Cygnus Rift from Zhang et al., 2024) on the extinction jumps detected within 1.1 kpc. The colormap shows the distance of the largest extinction in the distance interval, and the thick black contours delineate the integrated extinction in [1, 2] mag. The dashed ellipse denote the “800 pc loop” identified in Zhang et al. (2024). The rectangles indicate areas for further distance measurements in the following.

### v2

@@ 5.1 All Sky Dust Map
This method does not require prior assumptions of the number of jump points. Indeed, due to the complex distribution of gas on the Galactic plane, it is difficult to preset how many extinction components there are along a certain LoS. For instance, the $D$-AA curve is not merely a simple step function, but involves combinations of different scenarios (see details in Appendix B). Our method can traverse the jump points along the LoS in a certain region. This can be applied to detect components with significant extinction across the full sky, effectively generating a 3D extinction map for whole sky (see Figures 4 and 5 for the sky map and top-down views).
@@
To map all LoS characteristics across the celestial sphere, we apply the cylindrical equal area rectangular projection, which projects the sphere onto a rectangular plane. We set the sampling size (represent spatial resolution) $L_{0}=1^{\circ}$ and sampling interval (grid size) $L=30^{\prime}$ for a guide map. Setting aside the notable variations in star density across high Galactic latitudes and the Galactic plane, we generate a uniformly sampled planar diagram with a $30^{\prime}$ grid spacing and $30^{\prime}$ overlap to enhance the spatial correlation. This standardized format facilitates easy comparison with other maps. However, unlike traditional extinction maps, the distribution of extinction jumps in our map is discrete in the dimension of distance. For the sky map, we perform integration across various distance slices, while for the face-on map, we employ a 2D histogram with extinction magnitude as the weighting factor.
@@ 5.1.1 A Dust Map based on SHEDR3
We utilized the dataset from SHEDR3 (by Mode I, $C=2$), providing measurements of extinction $A_{V}$ and distance, with the maximum star distance capped at 5 kpc. The integration of this star table with infrared data aids in detecting more distant components with deeper extinction. The magnitudes of all identified jump points adhere to an exponential distribution. To improve detection accuracy and focus on more significant jumps, we retain jumps with $\Delta A_{V}$ exceeding 0.15 mag 111 A sample of this catalog with the description of the columns can be found in Table 2. The full machine-readable table of extinction components is presented in 10.57760/sciencedb.24650, from which the 3D extinction maps and distances of MCs can be derived.. This threshold is set by considering three key reasons: (1) 0.15 mag represents a typical precision value for extinction in the $V$ band from SHEDR3 at $G=17$ mag (Anders et al., 2022); (2) the simulations in section 4.1 demonstrate the algorithm’s effectiveness in detecting jump points above this threshold (assuming $\sigma\sim 0.2-0.5$ mag); (3) The most important is that, at this threshold, molecular hydrogen might be experiencing a phase transition based on the conversion between the extinction magnitude and the total column density of all gas. Here, we apply the empirical relationship between HI and the reddening, given by $N_{\rm HI}/E(B-V)=8.8\times 10^{21}{\rm cm}^{-2}{\rm mag}^{-1}$ (Lenz et al., 2017), with an extinction coefficient $R_{V}=3.1$ (Cardelli et al., 1989; Schlafly & Finkbeiner, 2011). This implies that the extinction contributed by HI is $A_{V}({\rm HI})=0.35\times 10^{-21}N_{\rm HI}$. And the total extinction $A_{V}$ reflects a combination of atomic gas and molecular hydrogen, which is $N_{\rm H}=N({\rm HI})+2N({\rm H}_{2})\approx 1.9\times 10^{21}{\rm cm}^{-2}A_{V}$ (Bohlin et al., 1978). Based on empirical relationships, the column density range where molecular hydrogen appears is $log~N_{\rm H}({\rm cm}^{-2})\approx 20.1-20.8$ (Gillmon et al., 2006; Bellomi et al., 2020), corresponding to an extinction value of $A_{V}({\rm mag})\approx 0.06-0.3$, with 0.15 mag being a mid-range value. Concurrently, Dame et al. (2001) suggests an extinction threshold for CO of $\gtrsim 0.4$ mag, indicating that some CO-dark molecular gas might appear below $\sim 0.4$ mag (Kalberla et al., 2020; Skalidis et al., 2024). We expect that a higher sensitivity survey may detect weak CO emission in some CO-dark regions.
@@
$\mu^{\mathrm{fore}}$: The mean extinction in the prejump distance bin.
$\mu^{\mathrm{back}}$: The mean extinction in the postjump distance bin.
$\sigma^{\mathrm{fore}}$: The standard deviation of extinction in the prejump distance bin.
$\sigma^{\mathrm{back}}$: The standard deviation of extinction in the postjump distance bin.
@@
In Figure 4, the Mollweide projection plot of the sky map illustrates the distances associated with the most significant extinction jumps across all LoS. The diagram shows some notable nearby regions that likely match those identified MCs by the single-extinction-jump model (Yan et al., 2019a; Zucker et al., 2019). The advantages of our model are that (1) the model with multi-extinction-jump can well trace all extinction structures with considerable $\Delta A_{\lambda}$ and (2) all the detected dust components are unaffected by different modeling process, i.e., the “ramps” in the $D$-AA map (Zucker et al., 2019) and the overlapping between multiple gas layers in distance measurements (i.e., multiple jump features along LoS, see Zhang et al., 2024). For the majority of the LoS, particularly at high Galactic latitudes, MCs are usually isolated and correspond to the strongest extinction components, facilitating the identification of their associated extinction structures and distances (Sun et al., 2024a). In contrast, the Galactic plane often features multiple molecular gas layers associated with various extinction components, complicating the precise determination of their distances. Given the correlation between MC column density and extinction, we prioritize the analysis of more significant extinction components (see Sections 5.2.1 and 5.2.2).
@@
Recent studies utilizing Gaia parallaxes have shed new light on the structure of the solar neighborhood. For example, some nearest star-forming regions are found to be located near the surface of the Local bubble (e.g., Taurus, Ophiuchus, Lupus, Chamaeleon, and Corona Australis; see details in Zucker et al., 2022). Our results also show prominent extinction features toward these regions within 300 pc (see thick blue regions in Figure 4). Some intriguing extinction structures are also revealed by our map (see light yellow bands for kpc-scale linear structures in Figure 5), e.g., the spur-like structure between the Sagittarius and Local arms (“Split” in Lallement et al., 2019), the Radcliffe Wave (Alves et al., 2020), the Cepheus spur (Pantaleoni González et al., 2021), etc. Besides, the large cavity in $120^{\circ}\lesssim l\lesssim 135^{\circ}$ (green ellipse in Figure 5, Vergely et al., 2022) is also clear in our map. We also find more oval cavity structures that dust is scarce in $200^{\circ}\lesssim l\lesssim 210^{\circ}$, $280^{\circ}\lesssim l\lesssim 300^{\circ}$ (see red ellipses in Figure 5). We call them the Mon-Gem superbubble and the Carina superbubble, respectively. While conventional extinction mapping obscures these features due to severe LoS extinction and artificial structures, our analysis more clearly reveals their kpc-scale morphology. These giant bubbles are reminiscent of the vast cavities observed across NGC 628 (the Phantom Voids; Watkins et al., 2023; Barnes et al., 2023), likely reflecting the global impact of intense past star formation activity on the gas distribution and dynamics of the Galactic disk. Future multiwavelength (e.g., X-ray and infrared) observations will be essential for further verification and to confirm their formation mechanisms.
Figure 6: Face-on maps from different star catalogs/techniques. Panels (a) and (b) present Polar coordinate dust maps of ZGR23. Panel (a) is produced by our method, and panel (b) is generated using a simple spatial binning technique as outlined by Zhang et al. (2023). Panel (c) is reproduced in Cartesian coordinates from the map in Sect 5.1.1. Panel (d) is from Vergely et al. (2022), using a hierarchical-inversion technique and two merged star catalogs.
@@ 5.1.2 Qualitative Comparison of Different Dust Maps
Panels (a) and (b) in Figure 6 both utilize the ZGR23 catalog in polar coordinates, where (a) represents our proposed method and (b) uses a simple spatial binning-based mapping technique. Notably, our method effectively suppresses finger-like artifacts and recovers structures in distant and low-extinction regions. Moreover, the reconstruction of the Perseus Arm in the second and third Galactic quadrants is particularly robust. While both methods maintain consistency in large-scale structures near the Sun, our approach demonstrates significant improvements in resolving finer details in complex environments.
@@
Panel (c) in Figure 6 presents a Cartesian projections of Figure 5, while (d) display hierarchical-inversion extinction map from Vergely et al. (2022). Both methodologies share three key traits: they utilize model-independent approaches, rely on the Gaia EDR3 astrometric backbone, and incorporate extensive optical/infrared survey data. Their top-down views align well in the solar neighborhood and inner Galaxy, but notable discrepancies emerge in some specific features. In essence, panel (c) shows a cleaner visualization by effectively mitigating “fingers-of-God” artifacts. Panel (d), while offering higher resolution, displays discrepancies such as the Scutum-Centaurus Arm appearing more distant and dispersed, and MCs like the one at (1.7, -3) kpc appearing farther away. Indeed, panel (c) from our work exhibits significantly greater spiral arm coherence than panel (d), especially in regions toward both the inner ($1.6\pm 0.2$ kpc and $2.6\pm 0.2$ kpc toward the Galactic center; Nogueras-Lara et al., 2021) and outer Galaxy (e.g., Sun et al., 2024b).
@@
Panels (a) and (c) in Figure 6 utilize the same methodology but rely on different catalogs. Their face-on views reveal systematic differences in their extinction estimates. While panel (a) results in a more compact distribution of dust clouds with reduced granularity, panel (c) more effectively resolves faint structures and exhibits prominent continuity in spiral arm features. The most notable differences between the two panels occur within the inner Galaxy, particularly toward the Galactic center, where panel (c) displays fewer elongated patterns in the face-on map but a sharper spiral feature.
@@
Previous studies (e.g., Green et al., 2019; Edenhofer et al., 2024) have conducted comparative analyses of various 3D extinction maps. However, significant methodological differences in dust mapping techniques and extinction correction approaches hinder direct comparisons between these studies. To establish definitive conclusions, controlled analyses using identical methods and stellar catalogs are essential. While variations exist in distance coverage and spatial resolution across different maps, these factors primarily affect small-scale structures rather than the overall large-scale morphology of the solar neighborhood. In summary, the CU-JADE method enables detection of MC distances, even from weak extinction jumps, while simultaneously improving 3D extinction mapping through reduction of fingers-of-God artifacts.
@@ 5.2.1 Cepheus Region
The MCs in the Cepheus region with different distances were roughly divided into three groups based on previous studies (Grenier et al., 1989; Yonekura et al., 1997): the nearby (200 pc $\lesssim d\lesssim$ 400 pc; ${\rm-9~km~s}^{-1}\lesssim v\lesssim$ 9 ${\rm km~s}^{-1}$), the distant (600 pc $\lesssim d\lesssim$ 1000 pc; ${\rm-26~km~s}^{-1}\lesssim v\lesssim$ 2 ${\rm km~s}^{-1}$), and those in the Perseus Arm ($d\gtrsim$ 3 kpc; ${\rm-37~km~s}^{-1}\lesssim v\lesssim$ ${\rm-30~km~s}^{-1}$). The distant group might overlap with the Cepheus Flare at high Galactic latitudes ($b\sim 13^{\circ}$), leading to large uncertainty and debate over the distances in this area. Zucker et al. (2019) also measured a group of distances at $\sim 950$ pc and 350~350 pc toward the region using a single-extinction-jump model.
@@
Figure 7: CfA $\rm{}^{12}CO$ and extinction map toward the Cepheus region. Panel (a) depicts the extinction between 200 and 500 pc using a colormap, while contours show the smoothed integrated intensity of $\rm{}^{12}CO$ emission within the LSR velocity range of -6 to 15 $\rm km~s^{-1}$. Panel (b) is similar to panel (a), but focuses on the distance range in (500, 1500) pc and LSR velocities in (-20, -6) $\rm km~s^{-1}$. Panel (c) is similar to panel (a), but presents data for distances in (1500, 3000) pc and LSR velocity in (-45, -20) $\rm km~s^{-1}$. Red circles indicate regions where extinction jumps were detected in right panels (d), (e), and (f). Vertical lines represent the distances of detected extinction jumps with uncertainties in $D$-AA plots. Gray dots are raw stellar data points, binned every 50 pc with black lines. Blue text and horizontal lines denote the mean extinction value after accounting for extinction jumps. Orange curves are binned radial velocities of stars from the Gaia DR3.
@@
Combined with CO data, our results give a more comprehensive view on this area. In Figure 7, we presented the zoom-in region between $l\in[100^{\circ},120^{\circ}]$ and $b\in[0^{\circ},20^{\circ}]$ from the all-sky map in Figure 4. Generally, MCs in Cepheus exhibit good correlations between the velocity and the corresponding distance, as well as the total gas column density and extinction (see panels (a), (b), and (c) in Figure 7). Based on the dust map in the Figure 7 (b), we have well determined some extended extinction structures from high Galactic latitudes to Galactic plane (see structures at 5001500500-1500 pc). These extinction structures match well with the molecular gas in distant groups (black contours in panel (b) of Figure 7). This large-scale structure may extend to the junction of Cepheus and Cassiopeia (see Figure 4), where the starburst cluster M120.1+3.0 is located at $\sim 1$ kpc (Sun et al., 2023). For MCs in the most negative velocities ($v\lesssim-20$ ${\rm km~s}^{-1}$), the corresponding extinction structures are primarily located in the Galactic plane and extend to the $b\sim 5^{\circ}$ (see circle f in panel (c) of Figure 7).
@@
In this direction, the nearby MCs at 200–500 pc (contours and colormap in panel (a) of Figure 7) are widely and nonuniformly distributed throughout the whole region. Studies suggest a concentration of young stars aged 1–5 Myr at distances of 330–370 pc (Szilágyi et al., 2021), indicating that these young stellar groups may be associated with the surrounding molecular gas at this distance. The nearby MCs with ${\rm-6~km~s}^{-1}\lesssim v\lesssim$ 15 ${\rm km~s}^{-1}$ are roughly ring-shaped, while the distant MCs (500–1500 pc; see circle e in panel (b) of Figure 7) are interspersed among them. These distant MCs near 1 kpc cause more significant extinction at $l\sim 109^{\circ},b\sim 13^{\circ}$ than structures at $\sim 300$ pc (see panel (e) of Figure 7). Our results show that there are two or even more gas layers toward the area near $l\sim 107^{\circ}-111^{\circ}$and $b\sim 13^{\circ}$ in the Cepheus Flare. Indeed, the Cephues Flare region exhibits multiple star-formaion episodes (about 7–30 Myr; Wang et al., 2025) that appear to have significantly reshaped the gas structure and spatial distribution within 500 pc.
@@
The $D$-$A_{V}$ diagrams (by Mode II and CC=2) of three directions (labeled (d), (e), and (f) in Figure 7) are collected from star samples within the circles. Combined with the left side-by-side plots, we further confirmed the close physical connection between CO clouds with different LSR velocities and the corresponding extinction structures in gas layers at different distances. At the same time, we plot the stellar radial velocities (RVs) after binning along the distance. The RVs of the stars only roughly conforms to the trend of the rotation curve. However, we find that the dispersion and the uncertainty of the RVs of stars are very large (see orange lines in $D$-$A_{V}$ diagram of Figure 7). Particular caution should be taken when using stellar RV in small velocity ranges and distance intervals to study the kinematics of their structures. On the contrary, the RV of a young star might be useful for describing the kinematics of the young stellar objects because of their smaller kinematical ages (Zhu et al., 2024).
@@
Figure 8: Overlay of the identified MWISP $\rm{}^{13}CO$ cloud structures (white contours) within 1.1 kpc (clouds in the Cygnus Rift from Zhang et al., 2024) based on the extinction jumps detected within 1.1 kpc. The colormap shows the distance of the largest extinction in the distance interval, and the thick black contours delineate the integrated extinction in [1, 2] mag. The dashed ellipse denotes the “800 pc loop” identified in Zhang et al. (2024). The rectangles indicate areas for further distance measurements in the following.